OpenStudy (anonymous):

Determine the derivative of g(x) = √[x](x^3 - x)

OpenStudy (sgadi):

\[g(x)= x^{\frac{5}{2}}-x^{\frac{3}{2}}\] \[g'(x)= \frac{5}{2}x^{\frac{3}{2}}-\frac{3}{2}x^{\frac{1}{2}}\] \[g'(x)= \frac{3\sqrt{x^3}}{2}-\frac{3\sqrt{x}}{2}\]

OpenStudy (anonymous):

The answer at the back says this: [rootx]/2 (7x^2 - 3) I got what you got as well..

OpenStudy (anonymous):

Yes this question I need help